OpenStudy (anonymous):

Kaitlyn enters a triathlon at school. In the triathlon she must swim, ride, and run. She must swim for x metres. She has to run 3 times as far as she swims, and ride 5 metres more than she swims. If the total length of the triathlon is 1000 m, how far must she ride?

OpenStudy (anonymous):

Where x is the length Kaitlyn swims: 1000 = x+3x+(x+5) 1000 = 5x+5 5x = 995 x = 995/5 = 199 Riding is 5m more than 5 Riding = 199+5 = 204